OpenStudy (anonymous):

Solve and Check. 3^2x-1 = 3^x+2

OpenStudy (zzr0ck3r):

\(3^{2x-1}=3^{x+2}\) take log base 3 of both sides \(\log_3(3^{2x-1})=\log_3(3^{x+2})\implies2x-1=x+2\) solve for x

OpenStudy (anonymous):

Oh, ok same bases now solve the exponents. I got this far. I'm a little confused about having x on both sides...do I bring x to one side?

OpenStudy (zzr0ck3r):

yes \(2x-1=x+2\\2x-1-x=2\\x-1=2\\x=3\)

OpenStudy (anonymous):

Ok, thank you so much! I got it now!